A nine-year, $270,000 promissory note bears interest at a rate of 8% compounded annually. What is its maturity value?
Answer #1

The formula to calculate Maturity Value (MV)

MV=P*(1+r)n

MV= Maturity Value

P= Principle Amount

r= rate of interest

n=number of compounding intervals between the time of deposit and maturity

Substituting with data given in the question;

P= 270,000

r= 8%

n= 9 years

thus

MV= 270,000*(1+8%)9

MV= 270,000*1.999004627

MV= 53,9731.25.